Soy bean meals is 16% protein, cornmeal is 8% protein. Howe many pounds of each should be mixed together in order to get 320 lbs. mixture that is 12% protein?
Answer by mananth(16946)

percent ---------------- quantity
Soyabean--16----------------x lb
cornmea--8----------------320-x lb
Total---12.00% ---------------- 320
16x+8(320-x)=12*320
16x+2560 -8x=3840
16x-8x=3840-2560
8x=1280
/8
x=160lb 16.00% soyabean
160lb 8.00% cornmeal
m.ananth@hotmail.ca